OpenStudy (anonymous):

8x^2+14x+5=0 completeing the squar

OpenStudy (solomonzelman):

\(\large\color{slate}{ 8x^2+14x+5=0 }\) \(\large\color{slate}{ 8x^2+14x=-5 }\) \(\large\color{slate}{ 8(x^2+\frac{14}{8}x)=-5 }\)

OpenStudy (solomonzelman):

Can you find the number that would make \(\large\color{blue}{ x^2+\frac{14}{8}x\color{red}{} }\) a perfect square trinomial once added?

OpenStudy (anonymous):

the anwser is 1.25 and .5 but i am not sure how to get them

OpenStudy (solomonzelman):

okay, can you answer the afore-posted question that I asked?

OpenStudy (anonymous):

14/8 1.75.. but i dont no what x is

OpenStudy (solomonzelman):

okay, lets first try to comprehend the concept of a perfect square trinomial. Do you know that: 1) \(\large\color{blue}{ x^2+4x+4 }\) 2) \(\large\color{blue}{ x^2-2x+1 }\) 3) \(\large\color{blue}{ x^2+6x+9 }\) 4) \(\large\color{blue}{ x^2+8x+16 }\) are perfect squares trinomials, and why this is so?

OpenStudy (solomonzelman):

(say no, if you don't... and don't be afraid of anything...)